A 100-volt electromotive force is applied to an RC series circuit in which the resistance is 200 ohms and the capacitance is 10−4 farad. Find the charge q(t) on the capacitor if q(0) = 0. Find the current i(t).



1
Expert's answer
2022-04-19T03:15:32-0400

Due to Kirchhoff's Law:

"U_R+U_C=E" (1)

where "U_R=iR=\frac{dq}{dt}R" , "U_C=\frac qC" , "E=100V".

Substitution "U_R" and "U_C" into (1) gives:

"\frac {dq}{dt}R+\frac qC=E"

"\frac {dq}{dt}+\frac q{RC}=\frac ER"

"\frac{1}{RC}=\frac{1}{200\cdot 10^{-4}}=50\Omega^{-1}F^{-1}"

"\frac ER=\frac{100}{200}=0.5\frac V\Omega"

"\frac {dq}{dt}+50q=0.5" (2)

The solutions to a nonhomogeneous equation are of the form

"q(t) = q_h(t) + q_p(t)" ,

where "q_h" is the general solution to the associated homogeneous equation and "q_p" is a particular solution.

The associated homogeneous equation:

"\frac {dq}{dt}+50q=0"

The general solution of this equation is determined by the roots of the characteristic equation:

"\lambda+50=0"

"\lambda=-50"

"q_h(t)=Ae^{-50t}" .

The particular solution of the differential equation:

"q_p(t)=B"

"q_p'(t)=0"

Substituting these into (2) we get:

"0+50B=0.5"

"B=0.01".

"q(t)=Ae^{-50t}+0.01"

Now we can apply initial condition to find A:

"q(0)=0"

"q(0)=A+0.01=0"

"A=-0.01"

"q(t)=0.01(1-e^{-50t})" ;

"i(t)=q'(t)=0.5e^{-50t}" .